# Large Language Models-guided Dynamic Adaptation for Temporal Knowledge Graph Reasoning

### v1

@@ Abstract
Temporal Knowledge Graph Reasoning (TKGR) is the process of utilizing temporal information to capture complex relations within a Temporal Knowledge Graph (TKG) to infer new knowledge. Conventional methods in TKGR typically depend on deep learning algorithms or temporal logical rules. However, deep learning-based TKGRs often lack interpretability, whereas rule-based TKGRs struggle to effectively learn temporal rules that capture temporal patterns. Recently, Large Language Models (LLMs) have demonstrated extensive knowledge and remarkable proficiency in temporal reasoning. Consequently, the employment of LLMs for Temporal Knowledge Graph Reasoning (TKGR) has sparked increasing interest among researchers. Nonetheless, LLMs are known to function as black boxes, making it challenging to comprehend their reasoning process. Additionally, due to the resource-intensive nature of fine-tuning, promptly updating LLMs to integrate evolving knowledge within TKGs for reasoning is impractical. To address these challenges, in this paper, we propose a Large Language Models-guided Dynamic Adaptation (LLM-DA) method for reasoning on TKGs. Specifically, LLM-DA harnesses the capabilities of LLMs to analyze historical data and extract temporal logical rules. These rules unveil temporal patterns and facilitate interpretable reasoning. To account for the evolving nature of TKGs, a dynamic adaptation strategy is proposed to update the LLM-generated rules with the latest events. This ensures that the extracted rules always incorporate the most recent knowledge and better generalize to the predictions on future events. Experimental results show that without the need of fine-tuning, LLM-DA significantly improves the accuracy of reasoning over several common datasets, providing a robust framework for TKGR tasks [^1] Code and data are available at: https://anonymous.4open.science/r/LLM-DA-1E6D.
@@ 1 Introduction
Temporal Knowledge Graphs (TKGs) [1, 2] are the structured representations of the real world, which incorporate the temporal dimension to analyze how relations between entities evolve over time. Temporal Knowledge Graph Reasoning (TKGR) focuses on leveraging historical information within TKGs to forecast future events. Prior research [3] on TKGR has primarily relied on temporal logical rules [4] or deep learning algorithms, such as graph neural networks [5] and reinforcement learning techniques [6]. However, the deep learning-based TKGRs often suffer from the lack of interpretability [7] and are difficult to dynamically update to accommodate new data in TKGs. While rule-based methods offer great interpretability and flexibility, effectively learning temporal logical rules and adapting them to new knowledge remains a huge challenge.
@@
Large Language Models (LLMs) [8], pretrained on large-scale text corpora, have exhibited extensive knowledge and reasoning ability. LLMs is able to effectively grasp intricate semantic and logical relationships within natural language, making them show remarkable performance across a wide range of tasks [9, 10, 11]. Recently, LLMs have also demonstrated surprising ability in temporal reasoning [12, 13, 14]. By utilizing their powerful contextual processing and pattern recognition abilities, LLMs can extract meaningful temporal patterns and complex temporal dependencies from historical data within TKGs [14, 15, 16], thereby significantly enhancing their temporal reasoning capabilities. Thus, leveraging the capabilities of LLMs [17] holds great promise for enhancing the performance of TKGR tasks.
@@ 2.1 Temporal Knowledge Graph Reasoning
Temporal Knowledge Graph Reasoning (TKGR) [25, 26] aims to leverage historical information within TKGs to forecast future events, which can be roughly categorized into two groups: Rules-based TKGR methods and Deep learning-based TKGR methods.
Rules-based TKGR methods enhance TKGs inference by leveraging temporal logical rules to accurately predict future events. TLmod [27] introduces a sophisticated pruning strategy to derive rules and selects the high-confidence rules for TKGR tasks. TLogic [4] learns temporal logical rules from TKGs based on temporal random walks, and subsequently feeds these rules into a symbolic reasoning module for predicting future events. TILP [28] proposes a differentiable framework for temporal logical rule learning, utilizing constrained random walks to enhance the learning process. TFLEX [29] advances beyond learning simplistic chain-like rules by proposing a temporal feature-logic embedding framework. Although temporal logical rules can reveal hidden temporal patterns within TKGR, effectively extracting these rules from TKG still remains a significant challenge.
Deep learning-based TKGR methods [30, 31] employ the deep learning techniques to capture the hidden temporal patterns to predict the future events in TKGs. RE-NET [32] uses a recurrent event encoder and a neighborhood aggregator to encode historical facts and model their connections, enhancing future event predictions in TKGs. Based on RE-NET, RE-GCN [33] further employs RGCN and GRU to aggregate neighboring messages and model the temporal dependency. CyGNet [34] employs a copy-generation mechanism for capturing global repetition frequencies. TiRGN [35] integrates both local and global historical data to capture the sequential, repetitive, and cyclical patterns inherent in historical data. However, the deep neural networks adopted by these methods often lack interpretability, making it difficult to verify the predictions.
@@ 4.1 Temporal Logical Rules Sampling
where $P_{xy}(t)$ is the probability of transitioning from node $e_{x}$ to node $e_{y}$ at time $t$, and $w(t)=\text{exp}(-\lambda(t-T))$ denotes an exponential decay function that weights the edges based on the time difference, $T$ is the current time, $t$ is the timestamp of the edge, and $\lambda$ is the decay rate parameter, which controls the weight given to more recent times. In this way, more recent times (i.e., $t$ closer to $T$) are assigned greater weight. The detailed theoretical analysis is shown in Appendix A. Through the constrained Markovian random walks on the historical data, LLM-DA extracts the temporal rules $\mathcal{S}$ from the sampled temporal paths.
@@ 4.2 Rule Generation
The detailed instruction for Rule Generation can be found in the Appendix B.1.
@@ 4.3 Dynamic Adaptation
Extracted Rules from Current Data. To address the issue of the broad range of rules generated by LLMs, LLM-DA explores constrained Markovian random walks to extract temporal logical rules from current data, which serves as a standard to constrain the scope of dynamic adaptation. Through iteratively invoking LLMs, the accuracy of LLMs in predicting future events can be enhanced. The detailed prompt for Dynamic Adaptation refers to Appendix B.2.
@@ 5.1 Experimental Settings
$\spadesuit$ RE-NET [32] 0.388 0.290 0.436 0.576 0.441 0.332 0.512 0.650
RE-GCN [33] 0.425 0.320 0.476 0.627 0.478 0.371 0.535 0.682
TiRGN [35] 0.441 0.341 0.497 0.650 0.495 0.389 0.559 0.703
TLogic [4] 0.390 0.295 0.437 0.573 0.459 0.360 0.518 0.646
$\clubsuit$ PPT [18] 0.384 0.289 0.425 0.570 0.389 0.286 0.434 0.586
Llama-2-7b-CoH [20] 0.349 0.470 0.591 0.386 0.541 0.699
Vicuna-7b-CoH [20] 0.328 0.457 0.656 0.392 0.546 0.707
GPT-NeoX [37] 0.334 0.460 0.565
Mixtral-8x7B-CoH [19] 0.439 0.331 0.496 0.649 0.497 0.380 0.564 0.713
@@ 5.3 Analysis of Dynamic Adaptation
RQ1: The ablation study on different data without dynamic adaptation aims to evaluate the impact of dynamic adaptation on performance. Dynamic adaptation typically employs the current data to update the rules generated by LLMs based on historical data. To demonstrate the superiority, we compare three variations, including “LLM-DA w H”, “LLM-DA w C” and “LLM-DA w H+C”, and the experimental results are shown in Table 2.
Specifically, the “LLM-DA w H+C” outperforms “LLM-DA w H” and “LLM-DA w C”, indicating that large-scale historical data can provide general knowledge, while current data offers relevant knowledge. The combination of both enhances the prediction of future events. Furthermore, compared to “LLM-DA w H+C”, LLM-DA shows a significant improvement. This phenomenon demonstrates that the dynamic adaptation strategy can effectively integrate historical data and current data, leveraging the current data to update the rules generated by LLMs on historical data.
@@
RQ3: To further verify whether the number of iterations of the dynamic adaptation module affects
@@
the performance, we conduct the different numbers of iterations experiment on both datasets. As shown in Figure 5, the MRR performance exhibits an increasing trend as the number of iterations increases. These observations indicate that the dynamic adaptation strategy can continuously update the LLMs-generated rules with the latest events through iterations, thereby better adapting to the dynamic changes of TKGs. This further demonstrates the effectiveness and necessity of the dynamic adaptation strategy in handling the evolving nature of TKGs. Moreover, we conduct the parameter analysis experiment to validate the impact of the weight $\alpha$ in Appendix C.7.
@@ 6 Conclusion
Appendix A Theoretical Analysis of the Constrained Markovian Random Walks
Appendix B Prompt for LLMs
B.1 Prompt for Rule Generation
@@
B.2 Prompt for Dynamic Adaptation

### v3

@@ Abstract
Temporal Knowledge Graph Reasoning (TKGR) is the process of utilizing temporal information to capture complex relations within a Temporal Knowledge Graph (TKG) to infer new knowledge. Conventional methods in TKGR typically depend on deep learning algorithms or temporal logical rules. However, deep learning-based TKGRs often lack interpretability, whereas rule-based TKGRs struggle to effectively learn temporal rules that capture temporal patterns. Recently, Large Language Models (LLMs) have demonstrated extensive knowledge and remarkable proficiency in temporal reasoning. Consequently, the employment of LLMs for Temporal Knowledge Graph Reasoning (TKGR) has sparked increasing interest among researchers. Nonetheless, LLMs are known to function as black boxes, making it challenging to comprehend their reasoning process. Additionally, due to the resource-intensive nature of fine-tuning, promptly updating LLMs to integrate evolving knowledge within TKGs for reasoning is impractical. To address these challenges, in this paper, we propose a Large Language Models-guided Dynamic Adaptation (LLM-DA) method for reasoning on TKGs. Specifically, LLM-DA harnesses the capabilities of LLMs to analyze historical data and extract temporal logical rules. These rules unveil temporal patterns and facilitate interpretable reasoning. To account for the evolving nature of TKGs, a dynamic adaptation strategy is proposed to update the LLM-generated rules with the latest events. This ensures that the extracted rules always incorporate the most recent knowledge and better generalize to the predictions on future events. Experimental results show that without the need of fine-tuning, LLM-DA significantly improves the accuracy of reasoning over several common datasets, providing a robust framework for TKGR tasks [^1] Code and data are available at: https://github.com/jiapuwang/LLM-DA.git.
@@ 1 Introduction
Temporal Knowledge Graphs (TKGs) [1, 2] are the structured representations of the real world, which incorporate the temporal dimension to analyze how relations between entities evolve over time. Temporal Knowledge Graph Reasoning (TKGR) focuses on leveraging historical information within TKGs to forecast future events. Prior research [3, 4] on TKGR has primarily relied on temporal logical rules [5] or deep learning algorithms, such as graph neural networks [6, 7, 8] and reinforcement learning techniques [9]. However, the deep learning-based TKGRs often suffer from the lack of interpretability [10] and are difficult to dynamically update to accommodate new data in TKGs. While rule-based methods offer great interpretability and flexibility, effectively learning temporal logical rules and adapting them to new knowledge remains a huge challenge.
@@
Large Language Models (LLMs) [11], pretrained on large-scale text corpora, have exhibited extensive knowledge and reasoning ability [12]. LLMs is able to effectively grasp intricate semantic and logical relationships within natural language, making them show remarkable performance across a wide range of tasks [13, 14, 15]. Recently, LLMs have also demonstrated surprising ability in temporal reasoning [16, 17, 18]. By utilizing their powerful contextual processing and pattern recognition abilities, LLMs can extract meaningful temporal patterns and complex temporal dependencies from historical data within TKGs [18, 19, 20], thereby significantly enhancing their temporal reasoning capabilities. Thus, leveraging the capabilities of LLMs [21] holds great promise for enhancing the performance of TKGR tasks.
@@ 2.1 Temporal Knowledge Graph Reasoning
Temporal Knowledge Graph Reasoning (TKGR) [29, 30, 31, 32, 33, 34, 35] aims to leverage historical information within TKGs to forecast future events, which can be roughly categorized into two groups: Rules-based TKGR methods and Deep learning-based TKGR methods.
Rules-based TKGR methods [36] enhance TKGs inference by leveraging temporal logical rules to accurately predict future events. TLmod [37] introduces a sophisticated pruning strategy to derive rules and selects the high-confidence rules for TKGR tasks. TLogic [5] learns temporal logical rules from TKGs based on temporal random walks, and subsequently feeds these rules into a symbolic reasoning module for predicting future events. TILP [38] proposes a differentiable framework for temporal logical rule learning, utilizing constrained random walks to enhance the learning process. TFLEX [39] advances beyond learning simplistic chain-like rules by proposing a temporal feature-logic embedding framework. Although temporal logical rules can reveal hidden temporal patterns within TKGR, effectively extracting these rules from TKG still remains a significant challenge.
Deep learning-based TKGR methods [40, 41, 42, 43] employ the deep learning techniques to capture the hidden temporal patterns to predict the future events in TKGs. RE-NET [44] uses a recurrent event encoder and a neighborhood aggregator to encode historical facts and model their connections, enhancing future event predictions in TKGs. Based on RE-NET, RE-GCN [45] further employs RGCN and GRU to aggregate neighboring messages and model the temporal dependency. CyGNet [46] employs a copy-generation mechanism for capturing global repetition frequencies. TiRGN [47] integrates both local and global historical data to capture the sequential, repetitive, and cyclical patterns inherent in historical data. However, the deep neural networks adopted by these methods often lack interpretability, making it difficult to verify the predictions.
@@ 4.1 Temporal Logical Rules Sampling
where $P_{xy}(t)$ is the probability of transitioning from node $e_{x}$ to node $e_{y}$ at time $t$, and $w(t)=\text{exp}(-\lambda(t-T))$ denotes an exponential decay function that weights the edges based on the time difference, $T$ is the current time, $t$ is the timestamp of the edge, and $\lambda$ is the decay rate parameter, which controls the weight given to more recent times. In this way, more recent times (i.e., $t$ closer to $T$) are assigned greater weight. The detailed theoretical analysis is shown in Appendix B. Through the constrained Markovian random walks on the historical data, LLM-DA extracts the temporal rules $\mathcal{S}$ from the sampled temporal paths.
@@ 4.2 Rule Generation
The detailed instruction for Rule Generation can be found in the Appendix A.1.
@@ 4.3 Dynamic Adaptation
Extracted Rules from Current Data. To address the issue of the broad range of rules generated by LLMs, LLM-DA explores constrained Markovian random walks to extract temporal logical rules from current data, which serves as a standard to constrain the scope of dynamic adaptation. Through iteratively invoking LLMs, the accuracy of LLMs in predicting future events can be enhanced. The detailed prompt for Dynamic Adaptation refers to Appendix A.2.
@@ 5.1 Experimental Settings
$\spadesuit$ RE-NET 0.388 0.290 0.436 0.576 0.441 0.332 0.512 0.650
RE-GCN 0.425 0.320 0.476 0.627 0.478 0.371 0.535 0.682
TiRGN 0.441 0.341 0.497 0.650 0.495 0.389 0.559 0.703
TLogic 0.390 0.295 0.437 0.573 0.459 0.360 0.518 0.646
$\clubsuit$ PPT 0.384 0.289 0.425 0.570 0.389 0.286 0.434 0.586
Llama-2-7b-CoH 0.349 0.470 0.591 0.386 0.541 0.699
Vicuna-7b-CoH 0.328 0.457 0.656 0.392 0.546 0.707
GPT-NeoX 0.334 0.460 0.565
Mixtral-8x7B-CoH 0.439 0.331 0.496 0.649 0.497 0.380 0.564 0.713
@@ 5.3 Analysis of Dynamic Adaptation
RQ1: The ablation study on different data without dynamic adaptation aims to evaluate the impact of dynamic adaptation on performance. Dynamic adaptation typically employs the current data to update the rules generated by LLMs based on historical data. To demonstrate the superiority, we compare three variations, including “LLM-DA w H”, “LLM-DA w C” and “LLM-DA w H+C”, and the experimental results are shown in Table 2. Specifically, the “LLM-DA w H+C” outperforms “LLM-DA w H” and “LLM-DA w C”, indicating that large-scale historical data can provide general knowledge, while current data offers relevant knowledge. The combination of both enhances the prediction of future events. Furthermore, compared to “LLM-DA w H+C”, LLM-DA shows a significant improvement. This phenomenon demonstrates that the dynamic adaptation strategy can effectively integrate historical data and current data, leveraging the current data to update the rules generated by LLMs on historical data.
@@
RQ3: To further verify whether the number of iterations of the dynamic adaptation module affects the performance, we conduct the different numbers of iterations experiment on both datasets. As shown in Figure 5, the MRR performance exhibits an increasing trend as the number of iterations
@@
increases. These observations indicate that the dynamic adaptation strategy can continuously update the LLMs-generated rules with the latest events through iterations, thereby better adapting to the dynamic changes of TKGs. This further demonstrates the effectiveness and necessity of the dynamic adaptation strategy in handling the evolving nature of TKGs. Moreover, we conduct the parameter analysis experiment to validate the impact of the weight $\alpha$ in Appendix C.7.
@@ 6 Conclusion
Appendix A Prompt for LLMs
A.1 Prompt for Rule Generation
@@
A.2 Prompt for Dynamic Adaptation
@@
Appendix B Theoretical Analysis of the Constrained Markovian Random Walks